John slid the cup from Mary to Bill.

CAUSE(John, GO(cup, [_{Path} FROM(Mary), TO(Mary)]))

[figure]

John John

slid the cup from Mary to Bill CAUSE(x, GO(cup, [_{Path} FROM(Mary), TO(Mary)]))

[figure]

slid

CAUSE(x, GO(y, [_{Path} u, v]))

[figure: the cup cup; _{\begin{matrix}the \\ ⊥ \end{matrix} {cup \atop cup}}]

from Mary FROM(Mary)

[figure]

from ; FROM(x) Mary Mary

to Bill TO(Bill)

[figure]

\begin{matrix}to \\ TO(x) \end{matrix} Bill Bill

[figure: \overline{X} x; XP; _{_{X_{SPEC}}  \overline{X}}; XP; _{\overline{X} _{X_{SPEC}}}; \overline{X}; _{\overline{X} Y}; \overline{X}; _{Y \overline{X}}]

(a)  (b)  (c)  (d)  (e)

[figure: \begin{matrix}ϕ \\ ϕ \end{matrix}; ϕ(ψ); _{ϕ(x) ψ}; ϕ(ψ); _{ψ ϕ(x)}; ϕ; _{ϕ ⊥}; ϕ; _{⊥ ϕ}]

(i)  (ii)  (iii)  (iv)  (v)

The cup slid from John to Mary. GO(cup, [_{Path} FROM(John), TO(Mary)])

IP

Fracture

[figure]

The cup cup NP

Fracture

[figure]

The ; ⊥ ; N_{SPEC} cup cup N

slid from John to Mary GO(x, [_{Path} FROM(John), TO(Mary)]) IP

Fracture

.. .

.. . throw:

EVENT

[figure]

ACT THING [(Bob)] THING [(box)] MANNER “throwing”

effect EVENT

[figure]

GO THING (box)

{PATH \atop [ ]}
[figure]

TO THING (Bill)

C^{2}
C^{1}
[figure]
I^{2}
[figure]
D_{1}^{2}
D^{1}
[figure]

D^{0} The N^{2} cup

I^{1}
[figure]
I^{0}
[figure]

V_{2}^{0} slide I^{0} -ed

V^{2}
[figure]
V^{2}
[figure] [figure: V^{2}; _{D_{1}^{2} t V^{1} V_{2}^{0} t}] P^{2}
P^{1}
[figure]

P^{0} from D^{2} John

P^{2}
P^{1}
[figure]

P^{0} to D^{2} Mary

\begin{matrix}C & 0 \\ ∅ \end{matrix}